A stone is dropped into a river from a bridge 47.8 m above the water. Another stone is thrown vertically down 1.60 s after the first is dropped. Both stones strike the water at the same time. What is the initial speed of the second stone?
Answer #1

Let us assume the downwards direction as positive.

Gravitational acceleration = g = 9.81 m/s2

Height of the bridge = H = 47.8 m

Initial velocity of the first stone = V1 = 0 m/s

Time taken by the first stone to hit the water = T1

H = V1T1 + gT12/2

47.8 = (0)T1 + (9.81)T12/2

T1 = 3.122 sec

Initial velocity of the second stone = V2

Time taken by the second stone to hit the water = T2

The second stone is thrown 1.6 sec after the first stone and both the stones hit the water at the same time.

T2 = T1 - 1.6

T2 = 3.122 - 1.6

T2 = 1.522 sec

H = V2T2 + gT22/2

47.8 = V2(1.522) + (9.81)(1.522)2/2

V2 = 23.94 m/s

Initial speed of the second stone = 23.94 m/s